OpenStudy (anonymous):

4^14 5^14 over (20^3)^4

OpenStudy (anonymous):

we know that (a^b)^c = a^bc right? and also know that... (ab)^c = (a^c)(b^c) so it becomes: (4^14) (5^14)/(4^12)( 5^12) which is : 16 * 25